# Scalable Link Prediction on Large-Scale Heterogeneous Graphs with Large Language Models

### 24 Jan 2024 v1

@@ Abstract
Exploring the application of large-scale language models to graph learning is a novel endeavor. However, the vast amount of information inherent in large graphs poses significant challenges to this process. This paper focuses on the link prediction task and introduces LPNL (Link Prediction via Natural Language), a framework based on a large language model designed for scalable link prediction on large-scale heterogeneous graphs. We design novel prompts for link prediction that articulate graph details in natural language. We propose a two-stage sampling pipeline to extract crucial information from large-scale heterogeneous graphs, and a divide-and-conquer strategy to control the input token count within predefined limits, addressing the challenge of overwhelming information. We fine-tune a T5 model based on our self-supervised learning designed for for link prediction. Extensive experiments on a large public heterogeneous graphs demonstrate that LPNL outperforms various advanced baselines, highlighting its remarkable performance in link prediction tasks on large-scale graphs.
@@ 1 Introduction
The emergence of large language models (LLMs) has elevated the field of natural language processing to unprecedented levels Devlin et al. (2018); Raffel et al. (2020); Brown et al. (2020). Recently, some research Fatemi et al. (2023); Ye et al. (2023) has explored the use of LLMs to address the learning on graphs. A popular paradigm of using LLMs on graph learning is to transform graph problems and structures into texts, and then feed the texts to LLMs to obtain the predictions. Figure 1 shows two pipelines for link prediction Daud et al. (2020); Lichtenwalter et al. (2010) based on Graph Neural Networks (GNNs) and LLMs. The former employs GNNs for graph learning, followed by a matcher to compute similarity to link predict Zhang and Chen (2018); Cai et al. (2021). Meanwhile, the latter generates textual prompts based on sampled subgraphs and utilizes LLMs for prediction.
In contrast to the conventional approach of GNNs that aggregate encoded features via message passing Hamilton et al. (2017); Abu-El-Haija et al. (2019), LLMs exhibit superior learning capabilities and enhanced flexibility Tamkin et al. (2021); Liang et al. (2022). However, it remains under explored on how to effectively learn from larger and more complex graphs with LLMs, which poses high challenges to capturing distant information and richer semantics. An example is the link prediction on large-scale graphs: as the number of candidate nodes or description nodes for prediction grows, so does the text fed into LLMs. In this case, extensive inputs become unfeasible due to token length limitations.
Figure 1: GNNs and LLMs based pipelines for link prediction
@@
In this work, we explore the scalable link prediction with large language models on large-scale heterogeneous graphs(LSHGs). LSHGs Shi et al. (2016); Wang et al. (2022) can be considered as networks of connections that resemble the real world more closely than conventional graph structures, such as social network Robins et al. (2007) or academic networks Kong et al. (2019). The key challenges can be succinctly described as follows: 1) how to fomulate the prompt template for scalable link prediction task. 2) how to find out crucial information on LSHGs by sampling, enabling LLMs to capture it within limited inputs. 3) how to address lengthy prompts generated by an excess of candidate nodes.
@@
To tackle the above challenges, we propose LPNL (Link Prediction via Natural Language), a large language model based framework for scalable link prediction on large-scale graphs. We design novel prompts for link prediction that articulating graph details in natural language. This involves establishing a selective query prompt template, furnishing a description of the link prediction task, and integrating heterogeneous information concerning the source and candidate nodes.
@@
In dealing with vast amounts of relevant graph information within LSHGs, LPNL’s focus lies on the crucial segments. This approach helps us avoid the interference of superfluous contextual information while ensuring compliance with specified token limitations. We design a two-stage sampling pipeline that utilizes normalized degree-based heterogeneous subgraph sampling and personalized pagerank-based ranking. This pipeline selects more crucial node information from the graph, ensuring that the model focuses more on more informative nodes.
@@
Figure 2: The framework of LPNL on heterogeneous graphs. For an input heterogeneous graph with link prediction tasks, perform two-stage sampling on the source node and each candidate node, resulting in the generation of corresponding anchor nodes. Generate prompts based on anchor nodes, input them into a large language model, reconstruct candidates based on the prediction results, create new prompts for another round of prediction, and repeat this divide-and-conquer process to obtain a unique link prediction result.
@@
With a large number of candidate nodes and only a single link prediction query, the token length constraints make it challenging to fully describe all candidate node information. For this situation, we utilize a divide-and-conquer method. We partition the original node set into multiple equally sized smaller sets, sequentially input them into the link prediction pipeline to obtain partial answers, and recursively reconstruct the candidate set to obtain the final answer.
@@
We conduct extensive experiments on the Open Academic Graph (OAG) Huang et al. (2020) and fine-tune the language models T5 Raffel et al. (2020) to serve as the backbone models on the OAG for LPNL. The results demonstrate that LPNL outperforms various enhanced GNNs-based baselines significantly in performance. Furthermore, through extensive experimentation, we discover that LPNL also exhibits remarkable few-shot capability. Unlike traditional models, LPNL’s training merely requires simple alignment formatting, enabling swift convergence in predictions and showcasing superior performance compared to conventional models. Additionally, experiments demonstrate the model’s robust knowledge transferability, maintaining consistent performance across various cross-domain tasks. This further underscores the immense potential of large language models in addressing graph learning.
@@ 2.1 Preliminary
Figure 3 showcases heterogeneous type nodes and their heterogeneous relationships sourced from OAG. Formally, a heterogeneous graph is denoted by $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{A},\mathcal{R}\}$, where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of nodes and edges (links), respectively. Each node $v\in\mathcal{V}$ and each link $e\in\mathcal{E}$ are associated with their mapping function $\phi(v):v\rightarrow\mathcal{A}$ and $\varphi(e):e\rightarrow\mathcal{R}$. $\mathcal{A}$ and $\mathcal{R}$ represent the sets of nodes and edge types.
Given a source node $s$ and a set of candidate nodes $\mathcal{C}=\{c_{1},c_{2},...,c_{n}\}$ satisfying a existed meta-relation $\langle\phi(s),\varphi(e),\phi(c_{i})\rangle$ where $e\in\mathcal{E}$ and $c_{i}\in\mathcal{C}$. Then a standard link prediction task on heterogeneous graphs is to predict a candidate node $c\in\mathcal{C}$ for a source node $s$ with the highest probability of $\langle s,e,c\rangle$.
Finally, let $\mathcal{G}_{sub}^{h}(v)=\{\mathcal{V}_{v}^{h},\mathcal{E}_{v}^{h},\mathcal{A}_{v}^{h},\mathcal{R}_{v}^{h}\}$ denote the $h$-hop ego-subgraph around $v$, consisting of $h$-hop neighbor nodes of $v$ and all interconnecting edges. We also denote $\mathcal{N}^{h}(v)$ as the set of all neighbor nodes on $\mathcal{G}_{sub}^{h}(v)$, which means $\mathcal{N}^{h}(v)=\{v^{\prime}|v^{\prime}\in\mathcal{V}_{v}^{h},v^{\prime}\neq v\}$. And $\mathcal{AN}^{h}_{k}(v)$ is denoted as the sequence of top-$k$ anchor nodes which is selected from $\mathcal{N}^{h}(v)$. Note that the above definitions are all heterogeneous.
@@ 2.2 Prompt Design for Link Prediction
In order to comprehensively represent the link prediction task along with the essential graph information, we meticulously design a uniform prompt template $\mathcal{T}(\cdot)$ for heterogeneous link prediction. Its fundamental mode involves a selective query, presenting both the link prediction problem description and information regarding the source node and candidate nodes. It prompts the large language models to identify the node most likely to be linked within the candidate set.
@@
Subsequently, given a source node $s$ and the set of candidate nodes $\mathcal{C}$, we formally obtain the link prediction prompt template as follows:
@@
where $\mathcal{R}$ is the relation type between the source node and candidate nodes and $n$ is the number of candidate nodes. And $q(\mathcal{R})$ represents a link prediction query, e.g., "which $\phi(c)$ is linked by $\phi(u)$?". Notably, in the above equation, the addition operators are redefined as the textual concatenation with separators. To better enable the large-scale model in distinguishing between various types of heterogeneous nodes, we additionally assign distinct type identifiers to the backend of each node, e.g. a paper node: <p>[PA]. We also show a more intuitive prompt template example in Figure 1.
Our designed prompts do not encompass the link information in the graph. Instead, we describe the key nodes in textual form based on their order of importance. This is because of the complexity of inter-node connections, which often result in redundant contexts Fatemi et al. (2023), making it challenging for the LLMs to comprehend. Consequently, LLMs might reducing the emphasis on the node features, which are pivotal for our tasks. Nonetheless, the links among heterogeneous nodes remain crucial as they reflect their relationships and node significance. To address this, we utilize the following two-stage sampling approach to prioritize critical nodes, thereby better encapsulating graph information.
@@ 2.3 Two-Stage Sampling
In the previous subsection, we designed the unified prompt template for link prediction. However, as graph data becomes more complex, resembling the real world, employing a single prompt engineering approach becomes challenging in addressing practical application problems.
Firstly, in large-scale heterogeneous graphs, if Eq.(1) attempts to represent the node $v$ information using its all $h$-hop neighbors, i.e. $k=\lvert\mathcal{V}_{v}^{h}\rvert$ formally, it will lead to an uncontrollable prompt length. Anothor issue arises due to substantial variations in the degrees of different node types. For example, the number of nodes in the $h$-hop subgraph around a paper-type node $\lvert\mathcal{N}^{h}($<p>)|$\rvert$) | is significantly smaller than that around a field-type node $\lvert\mathcal{N}^{h}($<f>)|$\rvert$) |. The two problems pose significant challenges to the input and contextual comprehension of large models.
In this paper, we provide a two-stage sampling pipeline. The first stage aims to sample subgraphs $\mathcal{G}_{sub}^{h}(v)$ based on normalized degree from large-scale heterogeneous graphs while mitigating sampling bias caused by heterogeneous types. And then we obtain the top-$k$ anchor nodes sequence $\mathcal{AN}^{h}_{k}(v)$ through the second stage sampling with personalized pagerank to generate $d(v)$ in Eq.(1).The further details are as follows.
Normalized Degree based Sampling Inspired by previous studies  Hu et al. (2020); Leskovec and Faloutsos (2006), we sample heterogeneous subgraphs based on normalized degree. Specifically, it specifies the sampling probability of each hop’s neighbors as their normalized degree. Normalized degree is defined as the current node’s degree normalized among all nodes of the same type in the same layer awaiting sampling. Therefore, for the $l$-th layer subgraph sampling around central node $s$, the sampling probability of node $v$ can be described as follows:
@@
where node $u$ represents the neighbor node at the $l$-th layer within the subgraph, satisfying $u\in\mathcal{V}_{s}^{h}-\mathcal{V}_{s}^{h-1}$ and $\phi(v)=\phi(u)$.
The normalized degree based sampling ensures that our first stage of sampling do not ignore the differences between node types so that it does not lead to bias against certain node types, e.g., nodes with higher degrees were not indiscriminately considered more important. Thus, a similar number of different types of nodes in the subgraphs was maintained, preserving richer semantic information. Furthermore, previous studies have demonstrated that leveraging up to 3-hop connectivity is effective for achieving excellent performance Kipf and Welling (2016); Veličković et al. (2017); Hamilton et al. (2017). However, extending the information beyond 3-hop generally has a marginal impact on improvement and, in some cases, may even result in negative effects Cai and Wang (2020); Zhang et al. (2021). Thus, we set the maximum value for multi-hop to 2-hop or 3-hop to implement the two-stage sampling approach.
@@
The two-stage sampling restricts the generated link prediction prompt length to suit large model inputs while maximizing the retention of crucial neighborhood information pertaining to the target node within the subgraphs. It also makes use of the structural information on the graph, so that the generated anchor nodes $\mathcal{AN}^{h}_{k}(s)$ can be seen as a hub converting the graph structure into textual descriptions. This enables our prompts generated in Sec. 2.2 to encompass not only node features but also implicit structural information.
@@
Table 1: Open Academic Graph (OAG) Statistics.
@@ 2.4 Divide-and-Conquer Prediction
While the previous subsection’s sampling pipeline addressed the potential issue of prompts length caused by Eq.(1), a careful observation of Eq.(2) reveals that an excessive number of candidate nodes in link prediction, denoted as $|\mathcal{C}|$, also makes the prompt length uncontrollable. Especially in large-scale heterogeneous graphs, the number of candidate nodes is high, making it difficult to describe all candidate nodes within a window-sized prompt. For instance, in a link prediction task with 100 candidate nodes, each node requires an average of approximately 250 tokens in the prompt for description. This would amount to a total of 25,000 tokens needed to describe all candidate nodes. This far exceeds the maximum token limit for a usual LLM’s context window. Furthermore, the excessive number of candidate nodes leads to redundant contexts, making it challenging for the LLMs to comprehend the input text.
@@
LPNL avoids the aforementioned token overload by employing a divide-and-conquer strategy Bentley (1980); Smith (1987). Firstly, we set a candidate length limit $\mathcal{L}$, ensuring that the length of all processed candidate sets does not exceed $\mathcal{L}$. Specifically, for an original candidate set $\mathcal{C}_{ori}$ of length $|\mathcal{C}_{ori}|$ where $|\mathcal{C}_{ori}|>\mathcal{L}$, we randomly divide it into $\lceil\frac{\mathcal{C}_{ori}}{\mathcal{L}}\rceil$ sub-candidate sets, i.e., $\mathcal{C}_{1}^{1},\mathcal{C}_{2}^{1},...,\mathcal{C}_{\lceil\frac{\mathcal{C}_{ori}}{\mathcal{L}}\rceil}^{1}$, satisfying $max(|\mathcal{C}_{1}^{1}|,|\mathcal{C}_{2}^{1}|,...,|\mathcal{C}_{\lceil\frac{\mathcal{C}_{ori}}{\mathcal{L}}\rceil}^{1}|)\leq\mathcal{L}$. As illustrated in Figure 1, by employing the fine-tuned large language model to predict the linked node with the maximum link probability for each sub-candidate set, we can subsequently eliminate low-probability candidate nodes. This process leads to more refined candidate sets by candidates reconstruction. And following this divide-and-conquer process by redividing candidate sets and making predictions, ultimately, we can obtain a unique prediction answer for the entire original candidate set $\mathcal{C}_{ori}$. Figure 4 provides a more intuitive illustration of the divide-and-conquer prediction, allowing us to observe the descent of candidate nodes and prompt tokens throughout the process.
Figure 4: Divide-and-conquer Prediction
2.5 Self-supervised Learning based Finetune

### 20 Feb 2024 v3

# LPNL: Scalable Link Prediction with Large Language Models

@@ Abstract
Exploring the application of large language models (LLMs) to graph learning is an emerging endeavor. However, the vast amount of information inherent in large graphs poses significant challenges to graph learning with LLMs. This work focuses on the link prediction task and introduces LPNL (Link Prediction via Natural Language), a framework based on large language models designed for scalable link prediction on large-scale heterogeneous graphs. We design novel prompts for link prediction that articulate graph details in natural language. We propose a two-stage sampling pipeline to extract crucial information from the graphs, and a divide-and-conquer strategy to control the input tokens within predefined limits, addressing the challenge of overwhelming information. We fine-tune a T5 model based on our self-supervised learning designed for link prediction. Extensive experimental results demonstrate that LPNL outperforms multiple advanced baselines in link prediction tasks on large-scale graphs.
@@ 1 Introduction
Heterogeneous graphs Shi et al. (2016) are commonly employed for modeling complex systems, wherein entities of diverse types interact with each other via various relations. Figure 1 shows the heterogeneous nodes and their relationships sourced from the Open Academic Graph (OAG) Huang et al. (2020). Link prediction Zhang and Chen (2018); Cai et al. (2021) is a fundamental task in graph learning. However, due to the vast quantity of nodes and edges with their complex structure, addressing the link prediction task on large-scale heterogeneous graphs is challenging.
Recently, some research Fatemi et al. (2023); Ye et al. (2023) has explored the use of large language models (LLMs) in graph learning. A popular paradigm of link prediction on graphs with LLMs is to transform graph problems and structures into description texts, and then feed the texts to LLMs to obtain the predictions. However, it remains under explored that how to perform scalable link prediction on large graphs through LLMs with the input window constraints, which poses serious challenges in capturing distant information and rich semantics. As the number of nodes increases, the text fed into LLMs grows. Consequently, extensive inputs become unfeasible due to token length limitations.
Figure 1: An example of heterogeneous graph
@@
In this work, we explore the scalable link prediction with large language models on large-scale heterogeneous graphs. The key challenges can be described as follows: 1) how to fomulate the prompt template for scalable link prediction task. 2) how to find out crucial information on large graphs, enabling LLMs to capture it within limited inputs. 3) how to address lengthy prompts generated by an excess of candidate neighbors. To tackle the above challenges, we propose LPNL (Link Prediction via Natural Language), a large language model based framework for scalable link prediction on large-scale graphs.
@@
We design novel prompts for link prediction that articulating graph details in natural language. This involves establishing a selective query prompt template, furnishing a description of the link prediction task, and integrating heterogeneous information concerning the source node and candidate neighbors.
@@
In dealing with vast amounts of relevant graph information within large graphs, LPNL selects crucial node information from the graph, ensuring that the model focuses more on them. We design a two-stage sampling pipeline that utilizes normalized degree-based heterogeneous subgraph sampling and personalized pagerank-based ranking. This approach avoids the interference of superfluous contextual information while ensuring compliance with specified token limitations.
@@
Figure 2: The framework of LPNL. For an input heterogeneous graph with link prediction tasks, LPNL consists of three steps: (1) conduct a two-stage sampling on the source node and each candidate neighbor from the original candidate set to acquire anchor nodes. (2) Generate prompts based on these anchor nodes and input them into LLMs for predictions. (3) Refine the candidate set based on prediction results and iteratively apply this divide-and-conquer process to obtain the distinct link prediction result $c^{*}$.
@@
With a large number of candidate neighbors, the token length constraints make it challenging to fully describe all candidate neighbor information. To address this issue, we employ a divide-and-conquer method. The original node set is partitioned into multiple sets with smaller size, which are sequentially input into the link prediction pipeline to obtain partial answers. Subsequently, we recursively refine the candidate set to predict the final answer.
@@
We conduct extensive experiments on the OAG and fine-tune the language model T5 Raffel et al. (2020) based on our self-supervised learning to serve as the backbone model for LPNL on the OAG . The results demonstrate that LPNL significantly outperforms various enhanced GNN-based baselines, achieving an average improvement of 30.52% on Hits@1. Furthermore, through extensive experimentation, LPNL also exhibits remarkable few-shot capability. Unlike traditional models training, LPNL’s fine-tuning merely requires simple alignment formatting, enabling swift convergence in predictions. Additionally, experiments demonstrate the model’s robust knowledge transferability, maintaining consistent performance across various cross-domain tasks. This further emphasizes that LPNL’s self-supervised fine-tuning is not confined to fixed graph labels, it can make direct predictions on different graphs without the need for additional learning.
@@ 2.1 Preliminary
Formally, a heterogeneous graph is denoted by $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{A},\mathcal{R}\}$, where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of nodes and edges (links), respectively. Each node $v\in\mathcal{V}$ and each link $e\in\mathcal{E}$ are associated with their mapping function $\phi(v):v\rightarrow\mathcal{A}$ and $\varphi(e):e\rightarrow\mathcal{R}$. $\mathcal{A}$ represents the set of node types, and $\mathcal{R}$ represents the set of edge types.
Given a source node $s$ and a set of candidate neighbors $\mathcal{C}=\{c_{1},c_{2},...,c_{n}\}$, satisfying a existed meta-relation $\langle\phi(s),\varphi(e),\phi(c_{i})\rangle$ where $e\in\mathcal{E}$ and $c_{i}\in\mathcal{C}$, a standard link prediction task on heterogeneous graphs aims to predict a candidate neighbor $c\in\mathcal{C}$ for a source node $s$ with the highest probability of $\langle s,e,c\rangle$.
Finally, let $\mathcal{G}_{sub}^{h}(v)=\{\mathcal{V}_{v}^{h},\mathcal{E}_{v}^{h},\mathcal{A}_{v}^{h},\mathcal{R}_{v}^{h}\}$ denote the $h$-hop ego-subgraph around $v$, consisting of $h$-hop neighbor nodes of $v$ and all interconnecting edges. We also denote $\mathcal{N}^{h}(v)$ as the set of all neighbor nodes on $\mathcal{G}_{sub}^{h}(v)$, which means $\mathcal{N}^{h}(v)=\{v^{\prime}|v^{\prime}\in\mathcal{V}_{v}^{h},v^{\prime}\neq v\}$. Additionally, $\mathcal{AN}^{h}_{k}(v)$ is denoted as the sequence of top-$k$ anchor nodes selected from $\mathcal{N}^{h}(v)$. Note that all the above definitions are heterogeneous.
@@ 2.2 Prompt Design for Link Prediction
In order to comprehensively represent the link prediction task along with the essential graph information, we meticulously design a uniform prompt template $\mathcal{T}(\cdot)$ for heterogeneous link prediction. Its fundamental mode involves a selective query, providing both the link prediction problem description and information regarding the source node and candidate neighbors. This prompts the large language models to identify the node most likely to be linked within the candidate set.
@@
Subsequently, given a source node $s$ and the set of candidate neighbors $\mathcal{C}$, we formally obtain the link prediction prompt template as follows:
@@
where $\mathcal{R}$ is the relation type between the source node and candidate neighbors and $n$ is the number of candidate neighbors. And $q(\mathcal{R})$ represents a link prediction query, e.g., "which $\phi(c)$ is linked by $\phi(u)$?". Notably, in the above equation, the addition operators are redefined as the textual concatenation with separators.
Author Disambiguation Example
[…]
Figure 3: The prompt example consists of three components: prefix_question: a selective question; source_node_description: the description of the source node and its corresponding anchor nodes; candidate_nodes_description: the description of candidate neighbors and the anchor nodes corresponding to each candidate neighbor.
To enhance the capability of the large language models in distinguishing between various types of heterogeneous nodes, we additionally assign distinct type identifiers to the backend of each node. For example, a paper node could be described as "<p>[PA]". Following the formal definition provided above, Figure 3 illustrates a more intuitive prompt example for author disambiguation.
Our designed prompts do not explicitly capture the link information between nodes in the graph. Instead, we choose to describe key nodes in textual form based on their order of importance. This decision arises from the complexity of inter-node connections, which often result in redundant contexts Fatemi et al. (2023), making it challenging for large language models to comprehend. Consequently, there is a risk of LLMs diminishing the emphasis on node features, which are pivotal for our tasks. Nonetheless, the links among heterogeneous nodes remain crucial as they reflect their relationships and node significance. In following Sec.2.3, we introduce a two-stage sampling approach to leverage structural information, prioritizing critical nodes and thereby enhancing the description of graph information.
@@ 2.3 Two-Stage Sampling
In the previous subsection, we designed the unified prompt template for link prediction. However, as graph data becomes more complex, resembling the real world, employing a single prompt engineering approach becomes challenging in addressing practical application problems. Firstly, in large-scale graphs, attempting to describe the node information of $v$ using all $h$-hop neighbors, i.e. $k=\lvert\mathcal{V}_{v}^{h}\rvert$ formally, as shown in Eq.(1), leads to an uncontrollable prompt length. Anothor issue arises due to substantial variations in the degrees of different node types. For example, the number of nodes in the $h$-hop subgraph around a paper node is significantly smaller than that around a field-type node. The two problems pose significant challenges to the input and contextual comprehension of LLMs.
In this work, we provide a two-stage sampling pipeline. The first stage aims to sample subgraphs $\mathcal{G}_{sub}^{h}(v)$ based on normalized degree from large-scale heterogeneous graphs while mitigating sampling bias caused by heterogeneous types. Subsequently, we obtain the top-$k$ anchor nodes sequence $\mathcal{AN}^{h}_{k}(v)$ through the second stage sampling with personalized pagerank to generate $d(v)$ in Eq.(1). The further details are as follows.
Normalized Degree based Sampling Inspired by previous studies  Hu et al. (2020); Leskovec and Faloutsos (2006), we adopt a strategy for sampling heterogeneous subgraphs based on normalized degree. Specifically, this approach specifies the sampling probability of each hop’s neighbors as their normalized degree. The normalized degree is defined as the node’s degree normalized among all nodes of the same type in the same layer. Therefore, for the $l$-th layer subgraph sampling around central node $s$, the sampling probability of node $v$ can be described as follows:
@@
where node $u$ represents the neighbor node at the $l$-th layer within the subgraph, satisfying $u\in\mathcal{V}_{s}^{h}\setminus\mathcal{V}_{s}^{h-1}$ and $\phi(v)=\phi(u)$.
The normalized degree based sampling in our first stage ensures that differences between node types are not ignored, preventing bias against certain node types (e.g., nodes with higher degrees are not indiscriminately considered more important). This approach maintains a similar number of different types of nodes in the subgraphs, thereby preserving richer semantic information. Furthermore, previous studies have demonstrated that leveraging up to 3-hop connectivity is effective for achieving excellent performance Kipf and Welling (2016); Veličković et al. (2017); Hamilton et al. (2017). However, extending the information beyond 3-hop generally has a marginal impact on improvement and, in some cases, may even result in negative effects Cai and Wang (2020); Zhang et al. (2021). Therefore, we set the maximum value for multi-hop to 2-hop or 3-hop in our two-stage sampling approach.
@@
The two-stage sampling restricts the generated link prediction prompt length to suit LLMs inputs while maximizing the retention of crucial neighborhood information pertaining to the target node within the subgraphs. It also makes use of the structural information on the graph, so that the generated anchor nodes $\mathcal{AN}^{h}_{k}(s)$ can be seen as a hub converting the graph structure into textual descriptions. This enables our prompts generated in Sec.2.2 to encompass not only node features but also implicit structural information.
@@
Table 1: OAG statistics.
Figure 4: For a link prediction task involving 100 candidate neighbors, we set the candidate length limit $L$ to 5. The candidate neighbors can be divided into 20 sets, followed by three rounds of divide-and-conquer. This process ultimately yields a unique prediction result.
@@ 2.4 Divide-and-Conquer Prediction
While the sampling pipeline addresses the potential issue of prompts length caused by Eq.(1), a careful observation of Eq.(2) reveals that an excessive number of candidate neighbors in link prediction, denoted as $|\mathcal{C}|$, also makes the prompt length uncontrollable. Especially in large-scale graphs, the high number of candidate neighbors poses a challenge in describing all of them within a single LLM’s input window. For instance, in a link prediction task with 100 candidate neighbors, each node requires an average of approximately 200 tokens in the prompt for description. This results in a total of 20,000 tokens needed to describe all candidate neighbors, far exceeding the maximum token limit for a usual LLM’s input window. Furthermore, the excessive number of candidate neighbors leads to redundant contexts, making it challenging for the LLMs to comprehend the input text.
@@
LPNL avoids the aforementioned token overload by employing a divide-and-conquer strategy. Figure 4 provides an intuitive example of the divide-and-conquer prediction, allowing us to observe the descent of candidate neighbors and prompt tokens throughout the process. We set a length limit $L$ for the candidate set, ensuring that the length of all processed candidate sets does not exceed $L$. We represent $\mathcal{C}^{i}_{j}$ as the $j$-th candidate set of the $i$-th divide-and-conquer round. Specifically, for an original candidate set of length $|\mathcal{C}^{0}|$ where $|\mathcal{C}^{0}|>L$, we randomly divide it into $m$ sub-candidate sets, ensuring $m=\lceil\frac{|\mathcal{C}^{0}|}{L}\rceil$. This results in sets denoted as $\mathcal{C}_{1}^{1},\mathcal{C}_{2}^{1},...,\mathcal{C}_{m}^{1}$, with the constraint that $max(|\mathcal{C}_{1}^{1}|,|\mathcal{C}_{2}^{1}|,...,|\mathcal{C}_{m}^{1}|)\leq L$.
As illustrated in Figure 1, by employing the fine-tuned large language models to predict the candidate neighbor of the source node with the maximum link probability for each sub-candidate set, we can subsequently eliminate low-probability candidate neighbors. And the process generates new candidate sets based on the predicted results by refining the candidate sets. Specifically, for the candidate sets $C_{j+1}^{i},C_{j+2}^{i},...,C_{j+k}^{i}$, a new candidate set $C_{k^{\prime}}^{i+1}$ is generated in the following round based on their prediction results. The values of $k$ and $k^{\prime}$ are determined based on the order of generation, ensuring that the condition $k\leq L$ is met. Following this divide-and-conquer process by refining candidate sets and making predictions, ultimately, we can obtain a unique prediction answer for the entire original candidate set $\mathcal{C}^{0}$.
2.5 Self-Supervised Fine-tuning